# Monte Carlo–Type Neural Operator for Differential Equations

### v1

@@ Abstract
The Monte Carlo-type Neural Operator (MCNO) introduces a framework for learning solution operators of partial differential equations (PDEs) by directly learning the kernel function and approximating the associated integral operator using a Monte Carlo-type approach. Unlike Fourier Neural Operators (FNOs), which rely on spectral representations and assume translation-invariant kernels, MCNO makes no such assumptions. The kernel is represented as a learnable tensor over sampled input–output pairs, and sampling is performed once, uniformly at random from a discretized grid. This design enables generalization across multiple grid resolutions without relying on fixed global basis functions or repeated sampling during training, while an interpolation step maps between arbitrary input and output grids to further enhance flexibility. Experiments on standard 1D PDE benchmarks show that MCNO achieves competitive accuracy with efficient computational cost. We also provide a theoretical analysis proving that the Monte Carlo estimator yields a bounded bias and variance under mild regularity assumptions. This result holds in any spatial dimension, suggesting that MCNO may extend naturally beyond one-dimensional problems. More broadly, this work explores how Monte Carlo-type integration can be incorporated into neural operator frameworks for continuous-domain PDEs, providing a theoretically supported alternative to spectral methods (such as FNO) and to graph-based Monte Carlo approaches (such as the Graph Kernel Neural Operator, GNO).
@@ I Introduction
Contributions. The key contributions of this paper are:
@@
  • We incorporate feature mixing into the Monte Carlo framework, capturing both spatial and cross-channel dependencies.
@@ II-A Operator Approximation
where $\kappa_{\phi}:\mathbb{R}^{2(d+d_{a})}\to\mathbb{R}^{d_{v}\times d_{v}}$ is a parameterized kernel function. This formulation extends the concept of neural networks to infinite-dimensional spaces, enabling the approximation of highly non-linear operators through the composition of linear integral transformations and non-linear activation functions.
@@ III MC-type Neural Operator Architecture
where $W$ is a linear transformation, $\sigma$ is a non-linear activation function, and $\hat{\mathcal{K}}_{\text{MC}}(a;\phi)v_{t}$ is the Monte Carlo-based kernel integral operator, learned by a direct representation using a weight tensor $\phi$. From now on, for ease of notation, we will denote the kernel integral and its MC estimator by $\mathcal{K}_{\phi}v_{t}$ and $\hat{\mathcal{K}}_{N,\phi}v_{t}$ respectively. The general architecture of MCNO is illustrated in Figure 1.
@@ III-A Monte Carlo Approximation for Integral Operators
where $\{y_{i}\}_{i=1}^{N}$ are points sampled uniformly at random from a discretized computational grid, with sampling performed once at the start of training.
Unlike the Fourier Neural Operator (FNO), which assumes translation invariance and parameterizes the kernel in Fourier space, MCNO learns $\kappa_{\phi}(x,y)$ directly in the spatial domain without structural constraints. This enables modeling of heterogeneous or non-stationary operators. The parameters $\phi$ are optimized end-to-end via backpropagation.
The fixed-sample strategy avoids repeated sampling during training, supports efficient batching, and integrates with structured-grid interpolation. The method is fully differentiable and lightweight to implement.
Algorithm 1 Monte Carlo-type Neural Operator (MCNO)
1:Input: Grid points, number of samples $N$, kernel parameters $\kappa_{\phi}$, and sampled points $\{y_{i}\}_{i=1}^{N}$.
2:Output: Updated representation $v_{t+1}(x)$.
3: Initialize the parameter $\phi$.
4: Estimate the kernel integral $(\mathcal{K}_{\phi}v_{t})(x)$ by :
$(\hat{\mathcal{K}}_{N,\phi}v_{t})(x)=\frac{1}{N}\sum_{i=1}^{N}\kappa_{\phi}(x,v_{t}(y_{i}))v_{t}(y_{i}).$
5: Interpolate onto the full grid
6: Update the representation:
$v_{t+1}(x)=\sigma\left(Wv_{t}(x)+(\hat{\mathcal{K}}_{N,\phi}v_{t})(x)\right).$
7: Optimize $\phi$ using the data loss function.
@@ III-B Efficient kernel interactions and feature mixing
MCNO leverages Einstein summation (via PyTorch’s einsum operation) to efficiently compute interactions between the learnable kernel and sampled input features, i.e., $\kappa_{\phi}(x,v_{t}(y_{i}))=\sum_{j}\phi_{j}\,v_{t,j}(y_{i})$, where $\phi$ is a learnable weight tensor. This avoids the overhead of explicit matrix multiplication and enables efficient parallelization on GPUs. Rather than aggregating over the entire grid, MCNO uses a subset of randomly sampled points to estimate the integral, maintaining efficiency while scaling to large problems.
Importantly, the learnable kernel operates over both spatial locations and feature channels. This means it not only aggregates spatial information from sampled points but also mixes input features across channels in a single operation. In contrast to models that handle spatial aggregation and channel mixing separately (e.g., with independent layers or shared weights), MCNO performs both jointly, enhancing expressiveness while keeping it computationally efficient.
@@ III-C Interpolation for grid reconstruction
Since the Monte Carlo integral is computed at a subset of sampled points, interpolation is required to reconstruct a structured grid representation compatible with standard neural architectures. In MCNO, this step maps sparse Monte Carlo estimates onto a dense discretized domain while preserving the local function structure.
MCNO employs continuous interpolation, typically using linear interpolation for smooth approximations while maintaining efficiency. However, for functions with sharp transitions, nearest-neighbor interpolation may prevent excessive smoothing. By decoupling function evaluation from grid constraints, MCNO enables flexible operator learning in both structured and irregular domains, adapting to complex spatial variations.
@@ III-D Linear Complexity
The Monte Carlo-type Neural Operator (MCNO) achieves linear computational complexity with respect to the number of samples $N$, offering an efficient method for approximating kernel integral operators. In contrast, the Fourier Neural Operator (FNO) computes the integral operator using the Fast Fourier Transform (FFT), which scales quasi-linearly as $\mathcal{O}(n\log n)$, where $n$ is the number of discretization points. While efficient in practice, the FFT introduces a logarithmic overhead that can affect scalability on large grids or in high-dimensional settings.
The kernel integral in MCNO is estimated by averaging over a fixed set of $N$ points sampled uniformly from a computational grid at the start of training. Under mild regularity assumptions on the integrand, we derive a theoretical bound on the approximation error, showing that both the bias and the variance of the estimator are controlled. Specifically, the bias decays with the resolution of the sampling grid $N_{\text{grid}}$, while the variance decreases with the number of samples $N$. This analysis confirms that MCNO maintains reliable approximation while keeping computational cost low.

### v2

@@ Abstract
The Monte Carlo-type Neural Operator (MCNO) introduces a framework for learning solution operators of one-dimensional partial differential equations (PDEs) by directly learning the kernel function and approximating the associated integral operator using a Monte Carlo-type approach. Unlike Fourier Neural Operators (FNOs), which rely on spectral representations and assume translation-invariant kernels, MCNO makes no such assumptions. The kernel is represented as a learnable tensor acting on the latent features at sampled points, with sampling performed once uniformly at random from the discretized grid. This design enables generalization across multiple grid resolutions without relying on fixed global basis functions or repeated sampling during training, while an interpolation step maps between arbitrary input and output grids to further enhance flexibility. Experiments on standard 1D PDE benchmarks show that MCNO achieves competitive accuracy with efficient computational cost. We also provide a theoretical analysis proving that the Monte Carlo estimator yields a bounded bias and variance under mild regularity assumptions. This result holds in any spatial dimension, suggesting that MCNO may extend naturally beyond one-dimensional problems. More broadly, this work explores how Monte Carlo-type integration can be incorporated into neural operator frameworks for continuous-domain PDEs, providing a theoretically supported alternative to spectral methods (such as FNO) and to graph-based Monte Carlo approaches (such as the Graph Kernel Neural Operator, GNO).
@@ I Introduction
Contributions. The key contributions of this paper:
@@
  • We integrate feature-channel mixing directly into the Monte Carlo aggregation, enabling the kernel to capture both spatial and cross-channel dependencies.
@@ II-A Operator Approximation
where $\kappa_{\phi}$ is a learnable tensor kernel that may depend on the spatial locations, input features, and the local latent representation $v_{t}(y)$.
This formulation extends the concept of neural networks to infinite-dimensional spaces, enabling the approximation of highly non-linear operators through the composition of linear integral transformations and non-linear activation functions.
@@ III MC-type Neural Operator Architecture
where $W$ is a linear transformation, $\sigma$ is a non-linear activation function, and $\hat{\mathcal{K}}_{\text{MC}}(a;\phi)v_{t}$ is the Monte Carlo-based kernel integral operator. From now on, for ease of notation, we will denote the kernel integral and its MC estimator by $\mathcal{K}_{\phi}v_{t}$ and $\hat{\mathcal{K}}_{N,\phi}v_{t}$ respectively. The general architecture of MCNO is illustrated in Figure 1.
@@ III-A Monte Carlo Approximation for Integral Operators
where $\{y_{i}\}_{i=1}^{N}$ are points sampled uniformly at random from a discretized computational grid, with sampling performed once at the start of training. This fixed-sample strategy avoids repeated sampling during training, supports efficient batching, and integrates with structured-grid interpolation.
Algorithm 1 MCNO — Forward Pass
1:Given: parameters $\phi$, $W$; samples $\{y_{i}\}_{i=1}^{N}$; grid; activation $\sigma$
2:Input: current representation $v_{t}(x)$
3:Output: $v_{t+1}(x)$
4: Estimate kernel:
$(\hat{\mathcal{K}}_{N,\phi}v_{t})(x)=\tfrac{1}{N}\sum_{i=1}^{N}\kappa_{\phi}\!\big(x,v_{t}(y_{i})\big)\,v_{t}(y_{i})$
5: Interpolate $(\hat{\mathcal{K}}_{N,\phi}v_{t})$ onto the full grid
6: Update: $v_{t+1}(x)=\sigma\!\left(Wv_{t}(x)+(\hat{\mathcal{K}}_{N,\phi}v_{t})(x)\right)$
@@ III-B Efficient kernel interactions and feature mixing
MCNO leverages Einstein summation (via PyTorch’s einsum operation) to efficiently compute interactions between the learnable kernel tensors $\phi=\{\phi_{i}\}_{i=1}^{N}$ and the sampled latent features, i.e., $\kappa_{\phi}(x,v_{t}(y_{i}))=\phi_{i}v_{t}(y_{i})$, where each $\phi_{i}\in\mathbb{R}^{d_{v}\times d_{v}}$ is a learnable tensor associated with sample index $i$.
Our implementation enables efficient batched computation and allows for GPU parallelization without requiring per-sample explicit matrix operations. Because each tensor $\phi_{i}$ acts directly on the feature channels, the kernel jointly aggregates information from the sampled spatial locations and mixes features across channels.
Unlike architectures that separate spatial aggregation from channel mixing, MCNO performs both within a unified operation, increasing expressiveness without adding computational overhead.
@@ III-C Interpolation for grid reconstruction
Since the Monte Carlo integral is computed at a subset of sampled points, interpolation is required to reconstruct a structured grid representation compatible with standard neural architectures. This step maps sparse Monte Carlo estimates onto a dense discretized domain while preserving the local function structure.
MCNO employs continuous interpolation, typically linear for smooth functions, though nearest-neighbor interpolation can be used when sharper transitions must be preserved. Because the reconstruction involves only local averaging and scales linearly with the number of grid points $N_{\text{grid}}$, it contributes modestly to the overall cost compared to the kernel evaluation.
@@ III-D Linear Complexity
MCNO achieves linear computational complexity with respect to the number of samples $N$, offering an efficient method for approximating kernel integral operators. In our implementation, the kernel takes the form $\kappa_{\phi}(x,v_{t}(y_{i}))=\phi_{i}v_{t}(y_{i})$, with each $\phi_{i}\in\mathbb{R}^{d_{v}\times d_{v}}$. This allows MCNO to compute kernel interactions only at the $N$ sampled points rather than across the full grid. This structure is a special case of the general kernel form used in our theoretical analysis. In contrast, the Fourier Neural Operator (FNO) computes the integral operator using the Fast Fourier Transform (FFT), which scales quasi-linearly as $\mathcal{O}(N_{\text{grid}}\log N_{\text{grid}})$, where $N_{\text{grid}}$ is the number of discretization points. While efficient in practice, the FFT introduces a logarithmic overhead that can become significant on large grids or in high-dimensional settings.
The Monte Carlo estimator further admits theoretical error guarantees. Under mild regularity assumptions, we show that the bias decreases with the grid resolution $N_{\text{grid}}$, while the variance decays with the number of samples $N$.